(Financial forecasting-percent of sales) Tulley Appliances, Inc. projects next year's sales to be $19.7 million. Current sales are at $14.8 million, based on current assets of $4.7 million and fixed assets of $4.8 on. The firm's net profit margin is 4.7 percent after axes. Tulley forecasts at current assets will rise in direct proportion the increase n sales, t fixed assets wil ncrease by ont 51 O. Currently T has $1.5 million in accounts payable (which vary directly with sales), S1.7 milion in long-term debt (due in 10 years), and common equity (including $3.9 million in retained earnings) totaling $6.6 million. Tulley plans to pay $460,000 in common stock dividends next year a. What are Tulley's total financing needs (that is, total assets) for the coming year? b. Given the firm's projections and dividend payment plans, what are its discretionary financing needs? c. Based on your projections, and assuming that the $103,000 expansion in fixed assets will occur, what is the largest increase in sales the firm can support without having to resort to the use of discretionary sources of financing? What are Tulley's total financing needs (that is, total assets) for the coming year? a. Estimate Tulley's financing needs by completing the pro forma balance sheet: (Round the percentages of sales to two decimal places and the balance sheet amounts to the nearest dollar.)
Pro Forma Balance Sheet Next Year % of Sales Current assets Net fixed assets Total assets Accounts payables Long-term debt Total liabilities Paid-in capital Retained earnings Common equity Total liabilities and common equity $
Answer #1

All financials below are in $ mn.

Current assets as % of sale = 4.7 / 14.8 = 31.76%

Accounts payables as % of sale = 1.5 / 14.8 = 10.14%

Please see the table below. In the last column i have explained how each value has been calculated.

Sales next year = S = 19.7

Addition to retained earnings = S x Profit margin - dividend payout = 19.7 x 4.7% - 0.460 = 0.4659.

While entering the figures, please round off the balance sheet figures to the nearest dollar.

Parameter Next year % of sales How it has been calculated?
Current Assets           6.26 31.76% =19.7 x 31.76%
Net fixed assets           4.90 = last year's value of 4.8 + 0.103
Total Assets         11.16 Sum of above two items
Accounts Payable           2.00 10.14% =19.7 x 10.14%
Long term debt           1.70 Same as last year
Total Liabilities           3.70 Sum of above two items
Paid in capital           2.70 Same as last year = 6.6 - 3.9
Retained earnings           4.37 Last year's balance, 3.9 + addition to retained earnings calculated above, 0.4659
Common equity           7.07 Sum of above two items
Total Liabilities & Equity         10.76 Total liabilities + Common equity

Part (a)

Total assets = 11.16

Part (b)

Discretionary financing needs = Total assets - total liabilities & equity without additional funding = 11.16 - 10.76 = 0.40 = $ 0.40 mn

Part (c)

If g is the growth rate then :

Current asset last year x g + Increase in net fixed assets = Accounts payables last year x g + S0 x (1 + g) x Profit margin - Dividend payout

Or, 4.7 x g + 0.103 = 1.5 x g + 14.8 x (1 + g) x 4.7% - 0.460

4.7g + 0.103 = 1.5g + 0.2356 + 0.6956g

Hence, g = (0.2356 - 0.103) / (4.7 - 1.5 - 0.6956) =  0.0529 = 5.29%

Hence, largest increase in sale = g = 5.29% = 5.29% x 14.8 =  0.7836 = $ 0.7836 mn